OpenStudy (anonymous):

@boyb39 help with a few question, i need to know if i'm right

OpenStudy (anonymous):

9.0?

OpenStudy (anonymous):

no x=11.8

OpenStudy (anonymous):

oh sorry this is the wrong question

OpenStudy (anonymous):

9.0?

OpenStudy (anonymous):

yes x=9.0298478316685 x=9.0

OpenStudy (anonymous):

I have a few more, so please bear with, and thank you very much

OpenStudy (anonymous):

44.8?

OpenStudy (anonymous):

yes the angel is 44.810922533528

OpenStudy (anonymous):

85?

OpenStudy (anonymous):

yes 84.7840914295459 =85

OpenStudy (anonymous):

3 more questions left to check,

OpenStudy (anonymous):

4.1?

OpenStudy (anonymous):

sorry i was doing other problem w=2.377

OpenStudy (anonymous):

no problem, don't be sorry, okay two more left

OpenStudy (anonymous):

1.8?

OpenStudy (anonymous):

1.1766709737697 i am not to sure about it

OpenStudy (anonymous):

i don't know the answer to this one

OpenStudy (anonymous):

it is like this x^2 =1^2+f^2-2*1*f cos(40)

OpenStudy (anonymous):

e^2 =1^2+f^2-2*1*f cos(40)

OpenStudy (anonymous):

OMG! thank you so much you don't know how much you've helped me, I finally can rest east and not fret over my work. You are truly an amazing human being for taking the time to help, even though you were not obliged to. I'm your BIGGEST fan.

OpenStudy (anonymous):

you are welcome :) glad i made you day

OpenStudy (anonymous):

have an amazing night, and god bless. :)